A flywheel in the form of a uniformly thick disk of radius 1.88 m, has a mass of 36.6 kg and spins counterclockwise at 371 rpm. Calculate the constant torque required to stop it in 2.75 min.
Answer #1
Concepts and reason

The required concepts to solve these questions are rotational kinematic equations, moment of inertia and torque.

Firstly, calculate the moment of inertia, calculate the angular acceleration by using rotational kinematic equations and then calculate the required torque to stop the car.

Fundamentals

The expression for angular velocity is expresses as follows,

ω=ω0+αt\omega = {\omega _0} + \alpha t

Here, α\alpha is the angular acceleration, tt is the time and ω0{\omega _0}is the initial angular velocity.

The expression for moment of inertia for disk is expresses as follows,

I=12mR2I = \frac{1}{2}m{R^2}

Here, RRis radius of the disk and mmis the mass of the flywheel.

The expression for torque is as follows,

τ=Iα\tau = I\alpha

Here, α\alpha is the angular acceleration and IIis the moment of inertia.

The expression for moment of inertia is,

I=12mR2I = \frac{1}{2}m{R^2}

Substitute 1.88m1.88{\rm{ m}}for RRand 36.6kg36.6{\rm{ kg}}for mmin the above equation.

I=12(36.6kg)(1.88m)2=64.67kgm2\begin{array}{c}\\I = \frac{1}{2}\left( {36.6{\rm{ kg}}} \right){\left( {1.88{\rm{ m}}} \right)^2}\\\\ = 64.67{\rm{ kg}} \cdot {{\rm{m}}^2}\\\end{array}

The expression for angular velocity is as follows,

ω=ω0+αt\omega = {\omega _0} + \alpha t …… (1)

Rearrange the equation for angular acceleration.

α=ωω0t\alpha = \frac{{\omega - {\omega _0}}}{t} …… (2)

Convert the unit of (ω0)\left( {{\omega _0}} \right)initial angular velocity (371rpm)\left( {371{\rm{ rpm}}} \right)torads1{\rm{rad}} \cdot {{\rm{s}}^{ - 1}}.

371rpm=(371rpm)(0.104rads11rpm)=38.58rads1\begin{array}{c}\\371{\rm{ rpm}} = \left( {371{\rm{ rpm}}} \right)\left( {\frac{{0.104{\rm{ rad}} \cdot {{\rm{s}}^{ - 1}}}}{{1{\rm{ rpm}}}}} \right)\\\\ = 38.58{\rm{ rad}} \cdot {{\rm{s}}^{ - 1}}\\\end{array}

Substitute 00 for ω\omega , 38.58rads138.58{\rm{ rad}} \cdot {{\rm{s}}^{ - 1}} for ω0{\omega _0} and2.75min2.75{\rm{ min}} for ttin the equation (2).

α=0(38.58rads1)(2.75min)(60s1min)=0.23rads2\begin{array}{c}\\\alpha = \frac{{0 - \left( {38.58{\rm{ rad}} \cdot {{\rm{s}}^{ - 1}}} \right)}}{{\left( {2.75{\rm{ min}}} \right)\left( {\frac{{60{\rm{ s}}}}{{1{\rm{ min}}}}} \right)}}\\\\ = - 0.23{\rm{ rad}} \cdot {{\rm{s}}^{ - 2}}\\\end{array}

The expression for torque is as follows,

τ=Iα\tau = I\alpha

Substitute 64.67kgm264.67{\rm{ kg}} \cdot {{\rm{m}}^2} for II , 0.23rads20.23{\rm{ rad}} \cdot {{\rm{s}}^{ - 2}} for α\alpha in the above equation of torque.

τ=(64.67kgm2)(0.23rads2)=14.87Nm\begin{array}{c}\\\tau = \left( {64.67{\rm{ kg}} \cdot {{\rm{m}}^2}} \right)\left( { - 0.23{\rm{ rad}} \cdot {{\rm{s}}^{ - 2}}} \right)\\\\ = - 14.87{\rm{ N}} \cdot {\rm{m}}\\\end{array}

The magnitude of the torque is,

τ=14.87Nm\tau = 14.87{\rm{ N}} \cdot {\rm{m}}

Ans:

The required torque to stop the wheel is 14.87Nm14.87{\rm{ N}} \cdot {\rm{m}}.